# Source reconstruction algorithms for coupled parabolic systems from internal measurements of one scalar state

### v3

@@ 6. Conclusions
Numerically, the inverse source problems were tackled using a descent method for an appropriate functional cost, including measures from the state and its first derivative in local subdomains (in line with the theoretical results). The parabolic systems were solved in 1D and 2D using finite element methods. We also presented a numerical analysis of the regularization term. We also compare the source reconstruction between local observations from one scalar component and all components of the state in every numerical experiment presented in this work.
Appendix A Proof of Theorem 2
In this appendix, we present the proof of Theorem 2, associated with Problem 1.
Proof.
Following the structure of the proof of Theorem 1, we divide this proof into three steps: spectral representation, null controllability, and Volterra equations.
Step 1. Spectral representation. First, we consider the representation of the solution of (3.2) in terms a series of the form (3.10). Here $Y_{k}(t)=(y^{k}_{1}(t),\dots,y^{k}_{n}(t))^{t}$ is the unique solution to the ordinary differential system
[…]
where $F_{k}=((f_{1},\varphi_{k})_{L^{2}(\Omega)},\dots,(f_{n},\varphi_{k})_{L^{2}(\Omega)})^{t}=:(f_{1}^{k},\dots,f_{n}^{k})^{t}$.
Now, by solving (A.1) we obtain
$\small Y_{k}(t)=\overbrace{\Biggl{(}\int\limits_{0}^{t}\tilde{\Phi}_{k}(t)\tilde{\Phi}_{k}^{-1}(s)\sigma(s)\,ds\Biggr{)}}^{M=(m_{ij}(t))}F_{k}=\Biggl{(}\sum\limits_{j=1}^{n}m_{1j}(t)f_{j}^{k},\,\,\sum\limits_{j=1}^{n}m_{2j}(t)f_{j}^{k},\,\,\dots,\sum\limits_{j=1}^{n}m_{nj}(t)f_{j}^{k}\Biggr{)}^{t}$} (A.2)
where $\tilde{\Phi}_{k}$ is a fundamental matrix associated with the ordinary differential system: $Z^{\prime}+\lambda_{k}DZ=0$ (see assumption H3).
Now, we define the sequence $\mathcal{B}:=\{\Psi^{0}_{k}\}_{k\in\mathbb{N}}$. In addition, we consider the integral representation (3.3), where now $W$ satisfies the system located in the right-hand side of (3.4). Thus, multiplying the identity (3.6) by the elements of $\mathcal{B}$ and integrating in space, we get
$\sigma(T)(F,\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}=\sigma(0)(W(T),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}+\int\limits_{0}^{T}\partial_{t}\sigma(T-s)(W(s),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}\,ds-(D\Delta Y(T),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}.$
Now, for every $k\in\mathbb{N}$, the last term in the above identity can be transformed as follows:
$-(D\Delta Y(T),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}=-(DY(T),\Delta\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}=\lambda_{k}\sum\limits_{j=1}^{n}\sum\limits_{i=1}^{n}d_{ij}y^{k}_{j}(T).$
Then, at this point, our reconstruction formula is given by
$(F,\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}=\frac{\sigma(0)}{\sigma(T)}(W(T),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}$
$+\frac{1}{\sigma(T)}\int\limits_{0}^{T}\partial_{t}\sigma(T-s)(W(s),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}\,ds+\frac{\lambda_{k}}{\sigma(T)}\sum\limits_{j=1}^{n}\sum\limits_{i=1}^{n}d_{ij}y^{k}_{j}(T),$
or equivalently (taking into account (A.2))
$\sum\limits_{j=1}^{n}\Biggl{(}1-\frac{\lambda_{k}}{\sigma(T)}\sum\limits_{\ell=1}^{n}\Biggl{(}\sum\limits_{i=1}^{n}d_{i\ell}\Biggr{)}m_{\ell j}(T)\Biggr{)}f_{j}^{k}=\frac{\sigma(0)}{\sigma(T)}(W(T),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}$ (A.3)
$\hskip 113.81102pt+\frac{1}{\sigma(T)}\int\limits_{0}^{T}\partial_{t}\sigma(T-s)(W(s),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}\,ds.$
Step 2. Controllability. In order to transform the global terms in (A.3) by local terms in the subdomain $\mathcal{O}\times(0,T)$, we apply the null controllability property for parabolic systems with coupling in the main operator (see Lemma 3). Then, denoting the adjoint state of $Y$ by $\Psi$, Lemma 3 allows to obtain a control function $U=U(\Psi_{k}^{0})\in L^{2}(0,s;L^{2}(\mathcal{O})^{n})$ such that the system
$\left\{\begin{array}[]{llll}-\partial_{t}\Psi-D^{*}\Delta\Psi=1_{\mathcal{O}}BU,&\mbox{in}&\Omega\times(0,s),\\ \Psi=0,&\mbox{on}&\partial\Omega\times(0,s),\\ \Psi(\,\cdot\,,s)=\Psi^{0}_{k}(\cdot),&\mbox{in}&\Omega,\end{array}\right$.\, (A.4)
satisfies
$\Psi(x,0)=0,\quad\forall x\in\Omega.$ (A.5)
Thus, multiplying the second system of (3.4) by $\Psi$ solution of (A.4) and (A.5), and integrating by parts in $L^{2}(0,s;L^{2}(\Omega)^{n})$, we obtain (after extending $U$ by zero at $(s,T)$)
$(W(s),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}=-(W,1_{\mathcal{O}}BU)_{L^{2}(0,s;L^{2}(\Omega)^{n}))}=-(W,1_{\mathcal{O}}BU)_{L^{2}(0,T;L^{2}(\Omega)^{n}))}.$ (A.6)
It is worth mentioning that the above identity is possible because $D$ is a constant matrix and the operator $A=-\Delta$ is self-adjoint; otherwise, the analysis might be significantly more complex.
Step 3. Volterra equations. Following the proof of Theorem 1, for every $k\in\mathbb{N}$, Lemma 4 and Lemma 5 guarantees the existence of a unique function $\Theta=(\theta_{1},\theta_{2},\dots,\theta_{n})^{t}$ satisfying (see (2.13) and (2.17))
$\overline{K^{*}}(\Theta)(x,t)=1_{\mathcal{O}}BU(x,t)\quad\mbox{ and}\quad\Theta(x,s)=0,\quad\forall k\in\mathbb{N},$ (A.7)
where $\overline{K^{*}}(\Theta):=(K^{*}\theta_{1},K^{*}\theta_{2},\dots,K^{*}\theta_{n})^{t}$, and $K^{*}\theta_{j}$ is the adjoint operator given in Lemma 5, for $j=1,\dots,n$. Then,
$(W(s),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}=-(W,B1_{\mathcal{O}}U)_{L^{2}(0,T;L^{2}(\Omega)^{n}))}=-(w_{n},K^{*}\theta_{n})_{H^{1}(0,T;L^{2}(\mathcal{O}))}.$ (A.8)
Finally, from (2.15) we have $y_{n}=Kw_{n}$, and in consequence
$(w_{n},K^{*}\theta_{n})_{H^{1}(0,T;L^{2}(\mathcal{O}))}=-(y_{n},\theta_{n})_{H^{1}(0,T;L^{2}(\mathcal{O}))}.$ (A.9)
Therefore, putting together (A.3), (A.8) and (A.9), we establish the formula (3.28), which completes the proof of Theorem 2. ∎
Appendix B Lipschitz stability via Carleman estimates
@@
$\left\{\begin{array}[]{llll}\partial_{t}Y-\Delta Y+QY=H,&\mbox{in}&\Omega\times(0,T),\\ Y=0,&\mbox{in}&\partial\Omega\times(0,T),\\ Y(\,\cdot\,,0)=0,&\mbox{in}&\Omega,\end{array}\right$. (B.1)
@@
Let $\mathcal{O}\subset\Omega\subset\mathbb{R}^{d}$ be a nonempty open subset and $r\leq 3$. Then, there exists a function $\alpha_{0}\in C^{2}(\overline{\Omega})$ (depending only on $\Omega$ and $\mathcal{O}$) and three positive constants $s_{1}=s_{1}(\Omega,\mathcal{O},T,\|q_{ij}\|_{\infty})$, $C=C(\Omega,\mathcal{O},T,\|q_{ij}\|_{\infty})$, and $\ell>3$ such that, for any $s\geq s_{1}$, the solution $Y$ of the problem (B.1) satisfies the inequality
$\begin{array}[]{lll}\sum\limits_{i=1}^{n}\mathcal{I}(3(n+1-i),y_{i})&\leq C\Biggl{(}s^{\ell}\displaystyle\iint\limits_{\mathcal{O}\times(0,T)}e^{-2s\alpha}\xi(t)^{\ell}|y_{n}|^{2}\,dx\,dt\\ &\hskip 28.45274pt+s^{r-3}\sum\limits_{i=1}^{n}\,\displaystyle\iint\limits_{\Omega\times(0,T)}e^{-2s\alpha}\xi(t)^{r-3}|h_{i}|^{2}\,dx\,dt\Biggr{)},\end{array}$ (B.2)
@@
$\mathcal{I}(r,z)=\iint\limits_{\Omega\times(0,T)}e^{-2s\alpha}\Biggl{(}(s\xi)^{r-4}|\partial_{t}z|^{2}+(s\xi)^{r-2}|\nabla z|^{2}+(s\xi)^{r}|z|^{2}\Biggr{)}\,dx\,dt,$ (B.3)
@@
The proof of Lemma 6 can be found in [29, Theorem 1.1]. We note that the first term of $\mathcal{I}(r,z)$, as well as the last term in the right-hand side of (B.2), do not appear explicitly in [29, Theorem 1.1]. However, by following the arguments therein, these additional terms can be incorporated without difficulty. Moreover, although Lemma 6 is valid for all $r\in\mathbb{R}$, we restrict ourselves to the case $r\leq 3$ in order to prove the Lipschitz stability result associated with our inverse source problem.
@@
Let $T^{\prime}=\frac{T}{2}$. Suppose that $\sigma\in W^{1,\infty}(0,T)$ and that there exists a positive constant $\gamma_{0}$ such that $|\sigma(T^{\prime})|\geq\gamma_{0}>0$. Assume further that $Y(\,\cdot\,,T^{\prime})=\tilde{Y}(\,\cdot\,,T^{\prime})$ in $\Omega$, where $Y$ and $\tilde{Y}$ denote solutions of (B.1) corresponding to the sources $\sigma(t)F(x)$ and $\sigma(t)\tilde{F}(x)$, respectively. Then, there exists a positive constant $C=C(\Omega,\mathcal{O},T^{\prime},T,\|q_{ij}\|_{\infty})$ such that the following estimate holds:
$\|F-\tilde{F}\|_{L^{2}(\Omega)^{n}}\leq C\|\partial_{t}y_{n}-\partial_{t}\tilde{y}_{n}\|_{L^{2}(0,T;L^{2}(\mathcal{O}))}.$ (B.4)
@@
Let $Y$ and $\tilde{Y}$ be solutions of (B.1) associated to the sources $\sigma F$ and $\sigma\tilde{F}$, respectively. By defining $U:=Y-\tilde{Y}$ and $Z:=\partial_{t}U$, we obtain the system:
$\left\{\begin{array}[]{llll}\partial_{t}Z-\Delta Z+QZ=(\partial_{t}\sigma)(F-\tilde{F}),&\mbox{in}&\Omega\times(0,T),\\ Z=0,&\mbox{on}&\partial\Omega\times(0,T).\end{array}\right$. (B.5)
Using that $\sigma\in W^{1,\infty}(0,T)$ and $|\sigma(T^{\prime})|\geq\gamma_{0}>0$, and by applying the Carleman estimate (B.2) to system (B.5), we obtain
$\begin{array}[]{lll}\displaystyle\sum\limits_{i=1}^{n}\mathcal{I}(3(n+1-i),z_{i})&\leq Cs^{\ell}\displaystyle\iint\limits_{\mathcal{O}\times(0,T)}e^{-2s\alpha}\xi(t)^{\ell}|z_{n}|^{2}\,dx\,dt\\ &\hskip 28.45274pt+Cs^{r-3}\displaystyle\sum\limits_{i=1}^{n}\,\displaystyle\iint\limits_{\Omega\times(0,T)}e^{-2s\alpha}\xi(t)^{r-3}|f_{i}-\tilde{f}_{i}|^{2}\,dx\,dt\end{array}$ (B.6)
@@
$\small{\begin{array}[]{lll}&s^{r-2}\displaystyle\int\limits_{\Omega}\xi^{r-3}(T^{\prime})|z_{i}(x,T^{\prime})|^{2}e^{-2s\alpha(x,T^{\prime})}\,dx\\ &=s\displaystyle\int\limits_{0}^{T^{\prime}}\frac{\partial}{\partial t}\Biggl{(}\int\limits_{\Omega}(s\xi)^{r-3}|z_{i}|^{2}e^{-2s\alpha}\,dx\Biggr{)}\,dt\\ &=s^{r-2}\displaystyle\int\limits_{\Omega}\int\limits_{0}^{T^{\prime}}\Biggl{(}(\partial_{t}\xi^{r-3})|z_{i}|^{2}+2\xi^{r-3}z_{i}\partial_{t}z_{i}-2s\xi^{r-3}(\partial_{t}\alpha)|z_{i}|^{2}\Biggr{)}e^{-2s\alpha}\,dx\,dt\\ &\leq C\displaystyle\iint\limits_{\Omega\times(0,T^{\prime})}\Biggl{(}(s\xi)^{r-2}|z_{i}|^{2}+(s\xi)^{r-4}|\partial_{t}z_{i}|^{2}+s^{r}\xi^{r-2}|z_{i}|^{2}+(s\xi)^{r-1}|z_{i}|^{2}\Biggr{)}e^{-2s\alpha}\,dx\,dt.\end{array}}$ (B.7)
@@ Acknowledgements
The first and third authors acknowledge the internal project 1109–11090202021 EAFIT University. Ignacio Brevis was supported by the Engineering and Physical Sciences Research Council (EPSRC), UK, under Grant EP/W010011/1.

### v4

@@ 6. Conclusions
From a numerical point of view, given the high computational cost of implementing reconstruction formulas via controllability and Volterra systems, we proposed a minimization-based approach. This method avoids repeatedly solving large-scale control problems and integral equations, while still preserving the core ideas of our theoretical framework. As such, it offers a more practical and scalable alternative for different space and time resolutions. Then, the inverse source problems were tackled using a descent method for an appropriate functional cost, including measures from the state and its first derivative in local subdomains (in line with the theoretical results). The parabolic systems were solved in 1D and 2D using finite element methods. We also presented a numerical analysis of the regularization term. We also compare the source reconstruction between local observations from one scalar component and all components of the state in every numerical experiment.
Appendix A Lipschitz stability via Carleman estimates
@@
$\left\{\begin{array}[]{llll}\partial_{t}Y-\Delta Y+QY=H,&\mbox{in}&\Omega\times(0,T),\\ Y=0,&\mbox{in}&\partial\Omega\times(0,T),\\ Y(\,\cdot\,,0)=0,&\mbox{in}&\Omega,\end{array}\right$. (A.1)
@@
Let $\mathcal{O}\subset\Omega\subset\mathbb{R}^{d}$ be a nonempty open subset and $r\leq 3$. Then, there exists a function $\alpha_{0}\in C^{2}(\overline{\Omega})$ (depending only on $\Omega$ and $\mathcal{O}$) and three positive constants $s_{1}=s_{1}(\Omega,\mathcal{O},T,\|q_{ij}\|_{\infty})$, $C=C(\Omega,\mathcal{O},T,\|q_{ij}\|_{\infty})$, and $\ell>3$ such that, for any $s\geq s_{1}$, the solution $Y$ of the problem (A.1) satisfies the inequality
$\begin{array}[]{lll}\sum\limits_{i=1}^{n}\mathcal{I}(3(n+1-i),y_{i})&\leq C\Biggl(s^{\ell}\displaystyle\iint\limits_{\mathcal{O}\times(0,T)}e^{-2s\alpha}\xi(t)^{\ell}|y_{n}|^{2}\,dx\,dt\\ &\hskip 28.45274pt+s^{r-3}\sum\limits_{i=1}^{n}\,\displaystyle\iint\limits_{\Omega\times(0,T)}e^{-2s\alpha}\xi(t)^{r-3}|h_{i}|^{2}\,dx\,dt\Biggr),\end{array}$ (A.2)
@@
$\mathcal{I}(r,z)=\iint\limits_{\Omega\times(0,T)}e^{-2s\alpha}\Biggl((s\xi)^{r-4}|\partial_{t}z|^{2}+(s\xi)^{r-2}|\nabla z|^{2}+(s\xi)^{r}|z|^{2}\Biggr)\,dx\,dt,$ (A.3)
@@
The proof of Lemma 5 can be found in [26, Theorem 1.1]. We note that the first term of $\mathcal{I}(r,z)$, as well as the last term in the right-hand side of (A.2), do not appear explicitly in [26, Theorem 1.1]. However, by following the arguments therein, these additional terms can be incorporated without difficulty. Moreover, although Lemma 5 is valid for all $r\in\mathbb{R}$, we restrict ourselves to the case $r\leq 3$ in order to prove the Lipschitz stability result associated with our inverse source problem.
@@
Let $T^{\prime}=\frac{T}{2}$. Suppose that $\sigma\in W^{1,\infty}(0,T)$ and that there exists a positive constant $\gamma_{0}$ such that $|\sigma(T^{\prime})|\geq\gamma_{0}>0$. Assume further that $Y(\,\cdot\,,T^{\prime})=\tilde{Y}(\,\cdot\,,T^{\prime})$ in $\Omega$, where $Y$ and $\tilde{Y}$ denote solutions of (A.1) corresponding to the sources $\sigma(t)F(x)$ and $\sigma(t)\tilde{F}(x)$, respectively. Then, there exists a positive constant $C=C(\Omega,\mathcal{O},T^{\prime},T,\|q_{ij}\|_{\infty})$ such that the following estimate holds:
$\|F-\tilde{F}\|_{L^{2}(\Omega)^{n}}\leq C\|\partial_{t}y_{n}-\partial_{t}\tilde{y}_{n}\|_{L^{2}(0,T;L^{2}(\mathcal{O}))}.$ (A.4)
@@
Let $Y$ and $\tilde{Y}$ be solutions of (A.1) associated to the sources $\sigma F$ and $\sigma\tilde{F}$, respectively. By defining $U:=Y-\tilde{Y}$ and $Z:=\partial_{t}U$, we obtain the system:
$\left\{\begin{array}[]{llll}\partial_{t}Z-\Delta Z+QZ=(\partial_{t}\sigma)(F-\tilde{F}),&\mbox{in}&\Omega\times(0,T),\\ Z=0,&\mbox{on}&\partial\Omega\times(0,T).\end{array}\right$. (A.5)
Using that $\sigma\in W^{1,\infty}(0,T)$ and $|\sigma(T^{\prime})|\geq\gamma_{0}>0$, and by applying the Carleman estimate (A.2) to system (A.5), we obtain
$\begin{array}[]{lll}\displaystyle\sum\limits_{i=1}^{n}\mathcal{I}(3(n+1-i),z_{i})&\leq Cs^{\ell}\displaystyle\iint\limits_{\mathcal{O}\times(0,T)}e^{-2s\alpha}\xi(t)^{\ell}|z_{n}|^{2}\,dx\,dt\\ &\hskip 28.45274pt+Cs^{r-3}\displaystyle\sum\limits_{i=1}^{n}\,\displaystyle\iint\limits_{\Omega\times(0,T)}e^{-2s\alpha}\xi(t)^{r-3}|f_{i}-\tilde{f}_{i}|^{2}\,dx\,dt\end{array}$ (A.6)
@@
$\small{\begin{array}[]{lll}&s^{r-2}\displaystyle\int\limits_{\Omega}\xi^{r-3}(T^{\prime})|z_{i}(x,T^{\prime})|^{2}e^{-2s\alpha(x,T^{\prime})}\,dx\\ &=s\displaystyle\int\limits_{0}^{T^{\prime}}\frac{\partial}{\partial t}\Biggl(\int\limits_{\Omega}(s\xi)^{r-3}|z_{i}|^{2}e^{-2s\alpha}\,dx\Biggr)\,dt\\ &=s^{r-2}\displaystyle\int\limits_{\Omega}\int\limits_{0}^{T^{\prime}}\Biggl((\partial_{t}\xi^{r-3})|z_{i}|^{2}+2\xi^{r-3}z_{i}\partial_{t}z_{i}-2s\xi^{r-3}(\partial_{t}\alpha)|z_{i}|^{2}\Biggr)e^{-2s\alpha}\,dx\,dt\\ &\leq C\displaystyle\iint\limits_{\Omega\times(0,T^{\prime})}\Biggl((s\xi)^{r-2}|z_{i}|^{2}+(s\xi)^{r-4}|\partial_{t}z_{i}|^{2}+s^{r}\xi^{r-2}|z_{i}|^{2}+(s\xi)^{r-1}|z_{i}|^{2}\Biggr)e^{-2s\alpha}\,dx\,dt.\end{array}}$ (A.7)
@@ Acknowledgements
The first and third authors acknowledge the internal project 1109–11090202021 EAFIT University. Ignacio Brevis was supported by the Engineering and Physical Sciences Research Council (EPSRC), UK, under Grant EP/W010011/1. This work was also supported by the London Mathematical Society (LMS), under the Scheme 5 Research Grant 52402.